40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 844     Entry time: Wed Jun 30 18:19:47 2010
Author: Jenna 
Type: Laser 
Category: GYRO 
Subject: Cavity Mirrors 

 Alastair and I went down to the lab today and changed the input mirror in the cavity to a two inch mirror and switched out the 6m mirrors with 9m ones. The input mirror has a transmission of 6515ppm, and the output mirror has a transmission of 5700ppm.

ELOG V3.1.3-